OpenStudy (anonymous):

A car is traveling on a highway. The distance (in miles) from its destination and the time (in hours) is given by the equation . What is the d-intercept of the line? (Remember the slope-intercept form is ) a. (0, 420) b. (420, 0) c. (0, 65) d. (65, 0)

OpenStudy (isaiah.feynman):

What??

OpenStudy (anonymous):

wooow

OpenStudy (isaiah.feynman):

Where is the equation?

OpenStudy (anonymous):

d=420-65t and y=mx+b

OpenStudy (isaiah.feynman):

So do you see the similarities of those two equations?

OpenStudy (anonymous):

noo

OpenStudy (isaiah.feynman):

I re arrange the first equation d=-65t+420 and y-mx+b. See anything?

OpenStudy (anonymous):

noo do you no the answer ??

OpenStudy (isaiah.feynman):

Yep. I want to guide you to it. Okay can you identify the terms in y=mx+b?

OpenStudy (anonymous):

can you just tell me the answer please

OpenStudy (isaiah.feynman):

Telling the answer never helps you, cos later if you see this type of question you won't be able to solve it still.